# Designing Phase Sensitive Probes of Monopole Superconducting Order

### v1

@@ Abstract
Distinct from familiar $s$-, $p$-, or $d$-wave pairings, the monopole superconducting order represents a novel class of pairing order arising from nontrivial monopole charge of the Cooper pair. In the weak-coupling regime, this order can emerge when pairing occurs between Fermi surfaces with different Chern numbers in, for example, doped Weyl semimetal systems. However, the phase of monopole pairing order is not well-defined over an entire Fermi surface, making it challenging to design experiments sensitive to both its symmetry and topology. To address this, we propose a scheme based on symmetry and topological principles to identify this elusive pairing order through a set of phase-sensitive Josephson experiments. By examining the discrepancy between global and local angular momentum of the pairing order, we can unveil the monopole charge of the pairing order. We demonstrate the proposed probe of monopole pairing order through analytic and numerical studies of Josephson coupling in models of monopole superconductor junctions. This work opens a promising avenue to uncover the unique topological properties of monopole pairing orders and to distinguish them from known pairing orders based on spherical harmonic symmetry.
@@ I Introduction
Monopole superconducting order [1] is an exotic family of three-dimensional topological superconducting orders. This pairing order arises from the nontrivial geometric phase of Cooper pairs [2] formed between two Fermi surfaces with different Chern numbers, which can occur in weak-coupling superconducting materials supporting Weyl Fermi surfaces [3, 4, 5, 6]. For example, when superconductivity is developed in a three-dimensional inversion symmetric magnetic Weyl system, intrinsically or by proximity, monopole harmonic pairing exists as the zero center-of-mass momentum pairing between two inversion-related Fermi surfaces enclosing Weyl points of opposite chiralities [1, 7, 8]. Unlike many other known topological superconductors, where only Bogoliubov-de Gennes (BdG) quasi-particle states cannot be adiabatically connected to a Bose-Einstein condensate [9, 10], the pairing order of a monopole superconductor also fails to be globally well-defined over the entire Fermi surface due to topological obstruction in the $\mathrm{U}(1)$ pairing phase. The topological obstruction in the pairing order also fundamentally changes the symmetry representation: monopole pairing order is no longer describable by familiar symmetry in terms of $s$-, $p$-, and $d$-waves based on spherical harmonics but is rather characterized by the so-called monopole harmonic functions [11, 12, 1, 7]. Moreover, the existence of pairing nodes in monopole superconducting order is independent of specific pairing mechanisms and is instead dictated by nonzero monopole charge of Cooper pairs, which results in unconventional pairing with nontrivial winding of pairing phase over an entire Fermi surface [2, 1].
@@
It has been known that, in the presence of a magnetic monopole in real space, the orbital angular momentum of a charged particle is no longer conserved. However, the total angular momentum, which includes the angular momentum of the electromagnetic field, remains conserved in both classical and quantum scatterings of charged particles [13, 14]. In a monopole superconductor, Cooper pairs scatter in the presence of a momentum space “magnetic” monopole. The total angular momentum of the pairing order, which we refer to as the “global angular momentum,” includes the contribution from the pair Berry flux and is a conserved physical observable that characterizes the global symmetry of the pairing order. However, even though the global angular momentum is conserved, the pairing order is not well-defined globally, i.e. for any choice of gauge, there exists at least a singular point at the Fermi surface at which the pairing phase cannot be well-defined. Only away from the singularity is the pairing order locally well-defined and able to be described in terms of spherical harmonic functions, which determine the “local angular momentum” of the pairing order. The shift, or say, the discrepancy, between the conserved global angular momentum and local angular momentum is a unique signature of the monopole charge of the pairing order.
@@
To detect superconducting pairing symmetries, phase-sensitive experiments have long played an essential role [15, 16, 17, 18, 19, 20, 21]. For example, the $d_{x^{2}-y^{2}}$ pairing symmetry of a high-temperature cuprate superconductor has been revealed by a corner Josephson junction. The relative sign difference of the pairing order along the $a$ and $b$ axes of a cuprate superconductor leads to a destructive Fraunhofer interference pattern at zero magnetic flux [20]. For $p$-wave topological superconductors, Majorana modes localized at boundaries have been proposed to contribute to single particle tunneling and the $4\pi$-periodic Josephson current phase relation [22, 23, 24, 25, 26, 27, 28, 29]. For monopole pairing order, the pairing phase is not well-defined over an entire Fermi surface, making it challenging to design experiments sensitive to both its symmetry and topology.
@@
This paper is organized as follows. In Sec. II, we discuss the first-order Josephson coupling between two uniform superconductors based on linear response theory, initially considering spin-orbit interactions at the interface and then extending to spin-orbit coupling in the bulk, emphasizing the role of the symmetry of superconducting orders in terms of a scattering form factor in this microscopic formulation. In Sec. III, we focus on Josephson junctions involving monopole superconductors, first deriving the form factor discussed in Sec. II based on linear response theory, and then analyzing the symmetry and topological principles which are independent of microscopic tunneling processes. We propose designing Josephson junctions aligned along or perpendicular to the high-symmetry rotational axis of the monopole superconductor to reveal the shift between the global and local angular momentum of the pairing order and extract the pair monopole charge. In Sec. IV, we design Josephson junctions to extract the global angular momentum of the monopole superconducting order, presenting results from both linear response theory and numerical results of a tight-binding model. Sec. V explores the design of Josephson junctions to probe the local angular momentum component by considering junctions between identical monopole superconductors. Finally, Sec. VI addresses the presence of spin-orbit interactions at the junction interface, proposing additional junction designs to distinguish monopole superconducting order from chiral spherical harmonic pairing orders.
II First order Josephson coupling in the presence of spin-orbit coupling
In this section, we employ linear response theory to derive the first order Josephson coupling [30, 31] in a superconductor-insulator-superconductor (SIS) junction in the presence of spin-orbit coupling in the bulk of the superconductors. Initially, we examine the scenario where only spin-orbit interactions occur at the junction interface, assuming the absence of spin-orbit coupling in the bulk. We emphasize the role of a form factor, which determines the symmetry selection rules for the first order Josephson current. Then, we additionally consider spin-orbit coupling in the bulk of superconductors. In the weak-coupling regime, by projecting to states near Fermi surfaces, we demonstrate that the bulk spin-orbit coupling can be treated as an effective interface spin-orbit interaction. Using this formalism, we can obtain the first order Josephson current in the presence of spin-orbit interactions in the bulk superconductors or at the interface.
@@ II First order Josephson coupling in the presence of spin-orbit coupling
${\sum_{\mathbf{k}}}^{\prime}\Psi^{\dagger}_{\alpha}(\mathbf{k})\left(\begin{array}[]{cc}\mathcal{H}_{\mathrm{kin},\alpha}(\mathbf{k})&\Delta_{\alpha}(\mathbf{k})e^{i\phi_{\alpha}}\\ \Delta^{\dagger}_{\alpha}(\mathbf{k})e^{-i\phi_{\alpha}}&-\mathcal{H}^{T}_{\mathrm{kin},\alpha}(-\mathbf{k})\end{array}\right)\Psi_{\alpha}(\mathbf{k}),$
where ${\sum_{\mathbf{k}}}^{\prime}$ denotes the sum over half the Brillouin zone and $\Psi_{\alpha}(\mathbf{k})=(c_{\alpha,\mathbf{k},\uparrow},c_{\alpha,\mathbf{k},\downarrow},c_{\alpha,-\mathbf{k},\uparrow}^{\dagger},c_{\alpha,-\mathbf{k},\downarrow}^{\dagger})^{\mathrm{T}}$ denotes the annihilation operator of Nambu spinor. Here, $c_{\alpha=L(R),\mathbf{k},\sigma}$ annihilates a single-particle state with momentum $\mathbf{k}$ and spin $\sigma=\uparrow,\downarrow$ at the left (right) side of the junction. $\mathcal{H}_{\mathrm{kin},\alpha}$ is the kinetic Hamiltonian kernel defined in the particle-hole channel, and $\Delta_{\alpha}e^{i\phi_{\alpha}}$ is the superconducting pairing matrix defined in the particle-particle channel with the overall $\mathrm{U}(1)$ pairing phase $e^{i\phi_{\alpha}}$ separated out here. $H_{\mathrm{link}}$ describes the single particle tunnelings across the insulating junction link and takes the form of the Bardeen-Josephson Hamiltonian [32, 33, 15] as follows,
@@
We now derive the first order Josephson current in this SIS junction at zero bias, where the dissipationless DC Josephson current running in thermal equilibrium is the only contributor to the tunneling current through the junction. Using linear response theory, to first order in the perturbation $H_{\mathrm{link}}$ in the expansion of the thermal weight $e^{-\beta H_{\mathrm{link}}}$ with $\beta=1/(k_{B}T)$, the DC Josephson current is given by
@@
where $\phi=\phi_{R}-\phi_{L}$ is the overall U($1$) phase difference of superconducting pairing across the junction, the trace is taken over spin space, and $p_{n}=(2n+1)\pi/\beta$ with $n\in\mathbb{Z}$ are fermion Matsubara frequencies. $\mathcal{F}_{\alpha}(\mathbf{k},ip_{n})$ is the anomalous Green’s function of the unperturbed BdG Hamiltonian $H_{\mathrm{BdG},\alpha}$ in Eq. (2). Generally, the total first order Josephson current can be expressed as
@@
in which $I_{c}$ is the Josephson critical current at zero bias.
@@
Here, the first term $e^{-i\phi}$ gives rise to $2\pi$-periodicity with respect to the pairing phase difference $\phi$ in the Josephson current $I_{J}(\phi)$. This reflects the first order process corresponding to a single Cooper pair tunneling through the junction barrier. The second term $w(E_{L;\mathbf{k}^{\prime}},E_{R,\mathbf{k}};\beta)\equiv\frac{1}{\beta}\sum_{ip_{n}}(p_{n}^{2}+E^{2}_{L,\mathbf{k}^{\prime}})^{-1}(p_{n}^{2}+E^{2}_{R,\mathbf{k}})^{-1}$ includes the temperature dependence and contribution from the unperturbed BdG quasiparticle energy dispersions, $E_{\alpha,\mathbf{k}}$. After frequency summation, it becomes
@@
Though the magnitude of the form factor depends on the microscopic details of tunneling at the junction interface via $T(\mathbf{k},\mathbf{k}^{\prime})$, whether it vanishes or not is sensitive to the superconducting orders at the two sides of the junction, $\Delta_{L}$ and $\Delta_{R}$, as well as the symmetry of the junction geometry. Therefore, the form factor $\mathfrak{F}(\mathbf{k},\mathbf{k}^{\prime})$ gives rise to symmetry selection rules which determine whether the first order Josephson current vanishes.
To demonstrate the symmetry selection rule provided by the form factor, consider a Josephson junction between a spin singlet and a spin triplet superconductor without spin-orbit coupling in the bulk superconductors but with spin-dependent tunneling at the junction interface. The pairing matrices of superconductors at two sides of the junction generally take the form $\Delta_{L}(\mathbf{k})=d_{L,0}(\mathbf{k})i\sigma_{y}$ and $\Delta_{R}(\mathbf{k})=(\mathbf{d}_{R}(\mathbf{k})\cdot\bm{\sigma})i\sigma_{y}$, with $d_{L,0}(\mathbf{k})$ the spatial part of the singlet pairing order and $\mathbf{d}_{R}(\mathbf{k})$ the $d$-vector of the triplet pairing order [34, 35]. Employing trace identities of Pauli matrices, we simplify the form factor for this singlet-triplet junction, denoted as $\mathfrak{F}_{\mathrm{sing}-\mathrm{trip}}$, to the following form,
@@
Consequently, in the absence of spin-orbit interaction at junction interface, $\mathbf{T}(\mathbf{k},\mathbf{k}^{\prime})=0$, the form factor vanishes and forbids first order Josephson coupling. This reflects the symmetry principle that when pairing orders are in different total spin channels, the first order Josephson coupling vanishes in the system where spin and orbital angular momentum are separately conserved.
Further consider Rashba-type spin-orbit interaction at the junction interface, as will be used later in Sec. VI, which has the form $\mathbf{T}(\mathbf{k},\mathbf{k}^{\prime})=T_{\mathrm{SO}}(\hat{\mathbf{n}}\times\hat{\mathbf{k}})\delta_{\mathbf{k},\mathbf{k}^{\prime}}$, with $\hat{\mathbf{n}}$ being the junction orientation. The form factor reduces to
@@
where we take $T_{0}(\mathbf{k},\mathbf{k}^{\prime})=T_{0}$, a constant, for simplicity. Together with Eq. (4), the first order Josephson current takes the form, $I_{J}=(8e/\hbar)T_{0}T_{\mathrm{SO}}\mathrm{Im}[e^{-i\phi}\sum_{\mathbf{k}}w(E_{L,\mathbf{k}},E_{R,\mathbf{k}};\beta)d_{L,0}(\mathbf{k})\hat{\mathbf{n}}\cdot(\hat{\mathbf{k}}\times\mathbf{d}_{R}^{*}(\mathbf{k}))]$, which agrees with the result obtained by Geshkenbein and Larkin [36]. Then, when $d_{L,0}(\mathbf{k}^{\prime})=\Delta_{L,0}$ for an $s$-wave superconductor, $\mathbf{d}_{R}(\mathbf{k})=\Delta_{R,0}(0,0,(k_{x}+ik_{y})/k_{F})$ for a chiral $p$-wave triplet superconductor with total spin $s_{z}=0$, and the junction is oriented along $\hat{\mathbf{n}}=\hat{z}\parallel\mathbf{d}_{R}$, the form factor in Eq. (10) vanishes, forbidding first order Josephson coupling even though there is nonvanishing Rashba spin-orbit interaction. This result can be understood from the following spin symmetry selection rules. In the spin-independent tunneling channel, as the singlet and triplet superconductors have different total spins, their first order Josephson coupling is zero. In the spin-dependent tunneling channel, the junction geometry determines the form of Rashba spin-orbit interaction as consisting of only in-plane spins which change $s_{z}$, the total spin-$z$ component of the Cooper pair, by $1$. However, $s_{z}$ of the pairing orders at two sides of the junction are identical, $s_{z}=0$. Hence, the first order Josephson coupling in this example vanishes.

### v2

@@ Abstract
Distinct from familiar $s$-, $p$-, or $d$-wave pairings, the monopole superconducting order represents a novel class of pairing order arising from nontrivial monopole charge of the Cooper pair. In the weak-coupling regime, this order can emerge when pairing occurs between Fermi surfaces with different Chern numbers in, for example, doped Weyl semimetal systems. However, the phase of monopole pairing order is not well-defined over an entire Fermi surface, making it challenging to design experiments sensitive to both its symmetry and topology. To address this, we propose a scheme based on symmetry and topological principles to identify this elusive pairing order through a set of phase-sensitive Josephson experiments. By examining the discrepancy between global and local angular momentum of the pairing order, we can unveil the monopole charge of the pairing order, including for models with pair monopole charge $|q_{p}|=1$, $2$, and $3$. We demonstrate the proposed probe of monopole pairing order through analytic and numerical studies of Josephson coupling in models of monopole superconductor junctions. This work opens a promising avenue to uncover the unique topological properties of monopole pairing orders and to distinguish them from known pairing orders based on spherical harmonic symmetry.
@@ I Introduction
Monopole superconducting order Li and Haldane (2018) is an exotic family of three-dimensional topological superconducting orders. This pairing order arises from the nontrivial geometric phase of Cooper pairs Murakami and Nagaosa (2003) formed between two Fermi surfaces with different Chern numbers, which can occur in weak-coupling superconducting materials supporting Weyl Fermi surfaces Armitage et al. (2018); Baidya and Vanderbilt (2020); Jeon et al. (2021); Xiao et al. (2022). For example, when superconductivity is developed in a three-dimensional inversion symmetric magnetic Weyl system, intrinsically or by proximity, monopole harmonic pairing exists as the zero center-of-mass momentum pairing between two inversion-related Fermi surfaces enclosing Weyl points of opposite chiralities Li and Haldane (2018); Li and Frazier (2024); Bobrow and Li . Unlike many other known topological superconductors, where only Bogoliubov-de Gennes (BdG) quasiparticle states cannot be adiabatically connected to a Bose-Einstein condensate Read and Green (2000); Volovik (2003), the pairing order of a monopole superconductor also fails to be globally well-defined over the entire Fermi surface due to topological obstruction in the $\mathrm{U}(1)$ pairing phase. The topological obstruction in the pairing order also fundamentally changes the symmetry representation: monopole pairing order is no longer describable by familiar symmetry in terms of $s$-, $p$-, and $d$-waves based on spherical harmonics but is rather characterized by the so-called monopole harmonic functions Wu and Yang (1976); Haldane (1983); Li and Haldane (2018); Li and Frazier (2024). Moreover, the existence of pairing nodes in monopole superconducting order is independent of specific pairing mechanisms and is instead dictated by nonzero monopole charge of Cooper pairs, which results in unconventional pairing with nontrivial winding of pairing phase over an entire Fermi surface Murakami and Nagaosa (2003); Li and Haldane (2018).
@@
It has been known that, in the presence of a magnetic monopole in real space, the orbital angular momentum of a charged particle is no longer conserved. However, the total angular momentum, which includes the angular momentum of the electromagnetic field, remains conserved in both classical and quantum scatterings of charged particles Goddard and Olive (1978); Atiyah and Hitchin (1985). In a monopole superconductor, Cooper pairs scatter in the presence of a momentum space “magnetic” monopole. The total angular momentum of the pairing order, which we refer to as the “global angular momentum,” includes the contribution from the pair Berry flux and is a conserved physical observable that characterizes the global symmetry of the pairing order. However, even though the global angular momentum is conserved, the pairing order is not well-defined globally, i.e. for any choice of gauge, there exists at least a singular point at the Fermi surface at which the pairing phase cannot be well-defined. Only away from the singular point is the pairing order locally well-defined and able to be described in terms of spherical harmonic functions, which determine the “local angular momentum” of the pairing order. The shift, or say, the discrepancy, between the conserved global angular momentum and local angular momentum is a unique signature of the monopole charge of the pairing order.
@@
To detect superconducting pairing symmetries, phase-sensitive experiments have long played an essential role Josephson (1962); Geshkenbein et al. (1987); Sigrist and M. Rice (1992); Wollman et al. (1993); Tsuei et al. (1994); Van Harlingen (1995); Tsuei and Kirtley (2000). For example, the $d_{x^{2}-y^{2}}$ pairing symmetry of a high-temperature cuprate superconductor has been revealed by a corner Josephson junction. The relative sign difference of the pairing order along the $a$ and $b$ axes of a cuprate superconductor leads to a destructive Fraunhofer interference pattern at zero magnetic flux Van Harlingen (1995). For $p$-wave topological superconductors, Majorana modes localized at boundaries have been proposed to contribute to single particle tunneling and the $4\pi$-periodic Josephson current phase relation Kitaev (2001); Kwon et al. (2004); Fu and Kane (2008); Chung and Zhang (2009); Sau et al. (2010); Alicea (2010); Grosfeld and Stern (2011); Alicea (2012). For monopole pairing order, the pairing phase is not well-defined over an entire Fermi surface, making it challenging to design experiments sensitive to both its symmetry and topology.
@@
This paper is organized as follows. In Sec. II, we discuss the first-order Josephson coupling between two uniform superconductors based on linear response theory, initially considering spin-orbit interactions at the interface and then extending to spin-orbit coupling in the bulk, emphasizing the role of the symmetry of superconducting orders in terms of a scattering form factor in this microscopic formulation. In Sec. III, we focus on Josephson junctions involving monopole superconductors, first deriving the form factor discussed in Sec. II based on linear response theory, and then analyzing the symmetry and topological principles which are independent of microscopic tunneling processes. We propose designing Josephson junctions aligned along or perpendicular to the high-symmetry rotational axis of the monopole superconductor to reveal the shift between the global and local angular momentum of the pairing order and extract the pair monopole charge. In Sec. IV, we design Josephson junctions to extract the global angular momentum of the monopole superconducting order, presenting results from both linear response theory and numerical results of a tight-binding model. Sec. V explores the design of Josephson junctions to probe the local angular momentum component by considering junctions between identical monopole superconductors. Finally, Sec. VII addresses the presence of spin-orbit interactions at the junction interface, proposing additional junction designs to distinguish monopole superconducting order from chiral spherical harmonic pairing orders.
II First-order Josephson coupling in the presence of spin-orbit coupling
In this section, we employ linear response theory to derive the first-order Josephson coupling Ambegaokar and Baratoff (1963); Sigrist and Ueda (1991) in a superconductor-insulator-superconductor (SIS) junction in the presence of spin-orbit coupling in the bulk of the superconductors. Initially, we examine the scenario where only spin-orbit interactions occur at the junction interface, assuming the absence of spin-orbit coupling in the bulk. We emphasize the role of a form factor, which determines the symmetry selection rules for the first-order Josephson current. Then, we additionally consider spin-orbit coupling in the bulk of superconductors. In the weak-coupling regime, by projecting to states near Fermi surfaces, we demonstrate that the bulk spin-orbit coupling can be treated as an effective interface spin-orbit interaction. Using this formalism, we can obtain the first-order Josephson current in the presence of spin-orbit interactions in the bulk superconductors or at the interface.
@@ II First-order Josephson coupling in the presence of spin-orbit coupling
${\sum_{\mathbf{k}}}^{\prime}\Psi^{\dagger}_{\alpha}(\mathbf{k})\left(\begin{array}[]{cc}\mathcal{H}_{\mathrm{kin},\alpha}(\mathbf{k})&\Delta_{\alpha}(\mathbf{k})e^{i\phi_{\alpha}}\\ \Delta^{\dagger}_{\alpha}(\mathbf{k})e^{-i\phi_{\alpha}}&-\mathcal{H}^{\mathrm{T}}_{\mathrm{kin},\alpha}(-\mathbf{k})\end{array}\right)\Psi_{\alpha}(\mathbf{k}),$
where ${\sum_{\mathbf{k}}}^{\prime}$ denotes the sum over half the Brillouin zone and $\Psi_{\alpha}(\mathbf{k})=(c_{\alpha,\mathbf{k},\uparrow},c_{\alpha,\mathbf{k},\downarrow},c_{\alpha,-\mathbf{k},\uparrow}^{\dagger},c_{\alpha,-\mathbf{k},\downarrow}^{\dagger})^{\mathrm{T}}$ denotes the annihilation operator of Nambu spinor. Here, $c_{\alpha=L(R),\mathbf{k},\sigma}$ annihilates a single-particle state with momentum $\mathbf{k}$ and spin $\sigma=\uparrow,\downarrow$ at the left (right) side of the junction. $\mathcal{H}_{\mathrm{kin},\alpha}$ is the kinetic Hamiltonian kernel defined in the particle-hole channel, and $\Delta_{\alpha}e^{i\phi_{\alpha}}$ is the superconducting pairing matrix defined in the particle-particle channel with the overall $\mathrm{U}(1)$ pairing phase $e^{i\phi_{\alpha}}$ separated out here. $H_{\mathrm{link}}$ describes the single particle tunnelings across the insulating junction link and takes the form of the Bardeen-Josephson Hamiltonian Bardeen (1961); Cohen et al. (1962); Josephson (1962) as follows,
@@
We now derive the first-order Josephson current in this SIS junction at zero bias, where the dissipationless DC Josephson current running in thermal equilibrium is the only contributor to the tunneling current through the junction. Using linear response theory, to first order in the perturbation $H_{\mathrm{link}}$ in the expansion of the thermal weight $e^{-\beta H_{\mathrm{link}}}$ with $\beta=1/(k_{B}T)$, the DC Josephson current is given by
@@
where $\phi=\phi_{R}-\phi_{L}$ is the overall U($1$) phase difference of superconducting pairing across the junction, the trace is taken over spin space, and $p_{n}=(2n+1)\pi/\beta$ with $n\in\mathbb{Z}$ are fermion Matsubara frequencies. $\mathcal{F}_{\alpha}(\mathbf{k},ip_{n})$ is the anomalous Green’s function of the unperturbed BdG Hamiltonian $H_{\mathrm{BdG},\alpha}$ in Eq. (2). Generally, the total first-order Josephson current can be expressed as
@@
where $I_{c}$ is the Josephson critical current at zero bias.
@@
Here, the first term $e^{-i\phi}$ gives rise to $2\pi$-periodicity with respect to the pairing phase difference $\phi$ in the Josephson current $I_{J}(\phi)$. This reflects the first-order process corresponding to a single Cooper pair tunneling through the junction barrier. The second term $w(E_{L;\mathbf{k}^{\prime}},E_{R,\mathbf{k}};\beta)\equiv\frac{1}{\beta}\sum_{ip_{n}}(p_{n}^{2}+E^{2}_{L,\mathbf{k}^{\prime}})^{-1}(p_{n}^{2}+E^{2}_{R,\mathbf{k}})^{-1}$ includes the temperature dependence and contribution from the unperturbed BdG quasiparticle energy dispersions, $E_{\alpha,\mathbf{k}}$. After frequency summation, it becomes
@@
Though the magnitude of the form factor depends on the microscopic details of tunneling at the junction interface via $T(\mathbf{k},\mathbf{k}^{\prime})$, whether it vanishes or not is sensitive to the superconducting orders at the two sides of the junction, $\Delta_{L}$ and $\Delta_{R}$, as well as the symmetry of the junction geometry. Therefore, the form factor $\mathfrak{F}(\mathbf{k},\mathbf{k}^{\prime})$ gives rise to symmetry selection rules which determine whether the first-order Josephson current vanishes.
To demonstrate the symmetry selection rule provided by the form factor, consider a Josephson junction between a spin singlet and a spin triplet superconductor without spin-orbit coupling in the bulk superconductors but with spin-dependent tunneling at the junction interface. The pairing matrices of superconductors at two sides of the junction generally take the form $\Delta_{L}(\mathbf{k})=d_{L,0}(\mathbf{k})i\sigma_{y}$ and $\Delta_{R}(\mathbf{k})=(\mathbf{d}_{R}(\mathbf{k})\cdot\bm{\sigma})i\sigma_{y}$, with $d_{L,0}(\mathbf{k})$ the spatial part of the singlet pairing order and $\mathbf{d}_{R}(\mathbf{k})$ the $d$-vector of the triplet pairing order Balian and Werthamer (1963); Leggett (1975). Employing trace identities of Pauli matrices, we simplify the form factor for this singlet-triplet junction, denoted as $\mathfrak{F}_{\mathrm{sing}-\mathrm{trip}}$, to the following form,
@@
Consequently, in the absence of spin-orbit interaction at junction interface, $\mathbf{T}(\mathbf{k},\mathbf{k}^{\prime})=0$, the form factor vanishes and forbids first-order Josephson coupling. This reflects the symmetry principle that when pairing orders are in different total spin channels, the first-order Josephson coupling vanishes in the system where spin and orbital angular momentum are separately conserved.
Further consider Rashba-type spin-orbit interaction at the junction interface, as will be used later in Sec. VII, which has the form $\mathbf{T}(\mathbf{k},\mathbf{k}^{\prime})=T_{\mathrm{SO}}(\hat{\mathbf{n}}\times\hat{\mathbf{k}})\delta_{\mathbf{k},\mathbf{k}^{\prime}}$, with $\hat{\mathbf{n}}$ being the junction orientation. The form factor reduces to
@@
where we take $T_{0}(\mathbf{k},\mathbf{k}^{\prime})=T_{0}$, a constant, for simplicity. Together with Eq. (4), the first-order Josephson current takes the form, $I_{J}=(8e/\hbar)T_{0}T_{\mathrm{SO}}\mathrm{Im}[e^{-i\phi}\sum_{\mathbf{k}}w(E_{L,\mathbf{k}},E_{R,\mathbf{k}};\beta)d_{L,0}(\mathbf{k})\hat{\mathbf{n}}\cdot(\hat{\mathbf{k}}\times\mathbf{d}_{R}^{*}(\mathbf{k}))]$, which agrees with the result obtained by Geshkenbein and Larkin Geshkenbein and Larkin (1986). Then, when $d_{L,0}(\mathbf{k}^{\prime})=\Delta_{L,0}$ for an $s$-wave superconductor, $\mathbf{d}_{R}(\mathbf{k})=\Delta_{R,0}(0,0,(k_{x}+ik_{y})/k_{F})$ for a chiral $p$-wave triplet superconductor with total spin $s_{z}=0$, and the junction is oriented along $\hat{\mathbf{n}}=\hat{z}\parallel\mathbf{d}_{R}$, the form factor in Eq. (10) vanishes, forbidding first-order Josephson coupling even though there is nonvanishing Rashba spin-orbit interaction. This result can be understood from the following spin symmetry selection rules. In the spin-independent tunneling channel, as the singlet and triplet superconductors have different total spins, their first-order Josephson coupling is zero. In the spin-dependent tunneling channel, the junction geometry determines the form of Rashba spin-orbit interaction as consisting of only in-plane spins which change $s_{z}$, the total spin-$z$ component of the Cooper pair, by $1$. However, $s_{z}$ of the pairing orders at two sides of the junction are identical, $s_{z}=0$. Hence, the first-order Josephson coupling in this example vanishes.